What does a control system do?

1. A control system maintains desired conditions in a physical system by adjusting selected variables in the system.

2. A feedback control system makes use of an output to influence an input to the same system.

Why is a control system necessary?

1. Safety

2. Environmental protection

3. Equipment protection

4. Smooth plant operation

5. Product quality

6. Profit optimization

7. Monitoring and diagnosis

How is control done?

1. Instrumentation for sensing

2. Computer calculations

3. Implementation of action computed by controller

What is necessary to design controllers?

1. Dynamic behavior of processes

2. Dynamic behavior of controllers

3. Interaction between dynamic systems

4. Engineering analysis

5. Common sense and engineering judgment

Course Overview

• Dynamic Process Modeling

• Dynamic Response

• Closed Loop Analysis

• Controller Tuning Guidelines

• Advanced Control Structures

• Interaction between Design and Control